Target, Inc. earned $2.00 per share five years ago. Its earnings this year were $3.20. What was the growth rate in earnings per share (EPS) over the 5-year period?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

3.2=2*(1+r/100)^5

(3.2/2)^(1/5)=(1+r/100)

(1+r/100)=1.0986

r=1.0986-1

=9.86%(Approx).